OpenStudy (anonymous):

Find a piecewise definition of f that does not involve the absolute value function f(x)=|x+10|

OpenStudy (anonymous):

i know it is x+10 and -x+10

OpenStudy (anonymous):

\[ |x+10| = \left\{\begin{array}{rcc} x + 10 & \text{if} & x \geq -10 \\ - x - 10& \text{if} & x < -10 \end{array} \right. \] would be one way

OpenStudy (anonymous):

it is not \(-x+10\) but rather \(-(x+10)=-x-10\) in the case where \(x+10<0\)